### v1

# Building Guardrails for Large Language Models

Abstract
As Large Language Models (LLMs) become more integrated into our daily lives, it is crucial to identify and mitigate their risks, especially when the risks can have profound impacts on human users and societies. Guardrails, which filter the inputs or outputs of LLMs, have emerged as a core safeguarding technology. This position paper takes a deep look at current open-source solutions (Llama Guard, Nvidia NeMo, Guardrails AI), and discusses the challenges and the road towards building more complete solutions. Drawing on robust evidence from previous research, we advocate for a systematic approach to construct guardrails for LLMs, based on comprehensive consideration of diverse contexts across various LLMs applications. We propose employing socio-technical methods through collaboration with a multi-disciplinary team to pinpoint precise technical requirements, exploring advanced neural-symbolic implementations to embrace the complexity of the requirements, and developing verification and testing to ensure the utmost quality of the final product.
Machine Learning, ICML
1 Introduction
Recent times have witnessed a notable increase in the utilization of Large Language Models (LLMs) like ChatGPT, attributed to their extensive and general capabilities (OpenAI, 2023). However, the rapid deployment and integration of LLMs have raised significant concerns regarding their risks including, but not limited to, ethical use, data biases, privacy and robustness (Huang et al., 2023b). In societal contexts, worries also include the potential misuse by malicious actors for activities such as spreading misinformation or aiding criminal activities, as indicated in studies by Kreps et al. (2022); Goldstein et al. (2023); Kang et al. (2023). In the scientific context, LLMs can be used in professional contexts, where there are dedicated ethical considerations and risks in scientific research (Birhane et al., 2023).
To address these issues, model developers have implemented a variety of safety protocols intended to confine the behaviors of these models to a more secure range of functions. The complexity of LLMs, characterized by intricate networks and numerous parameters, along with the closed-source nature (such as ChatGPT), present substantial hurdles. These complexities require different strategies compared to the pre-LLM era, which focus on white-box techniques, enhancing models by various regularisations and architecture adaptations during training. Therefore, in parallel to the reinforcement learning from human feedback (RLHF) and other training skills such as in-context training, the community moves towards employing black-box, post-hoc strategies, notably guardrails (Welbl et al., 2021; Gehman et al., 2020), which monitors and filters the inputs and outputs of trained LLMs. A guardrail is an algorithm that takes as input a set of objects (e.g., the input and/or the output of LLMs) and determines if and how some enforcement actions can be taken to reduce the risks embedded in the objects. For example, if an input to the LLMs is related to child exploitation, the guardrail may stop the input from being processed by the LLMs or adapt the output so that it becomes harmless (Perez et al., 2022). In other words, guardrails are to identify the potential misuse in the query stage and try to prevent the model from providing the answer that should not be given.
The difficulty in constructing guardrails often lies in establishing the requirements for them. E.g., AI regulations can be different across different countries, and in the context of a company, data privacy can be less serious than it is in the public domain. Nevertheless, a guardrail of LLMs may include requirements from one or more of the following categories: (i) Free from unintended responses e.g., offensive and hate speech (Section 3.1); (ii) Compliance to ethical principles such as fairness, privacy, and copyright (Section 3.23.3); (iii) Hallucinations and uncertainty (Section 3.4). In this paper, we do not include the typical requirement, i.e., accuracy, as they are benchmarks of the LLMs and arguably not the responsibilities of the guardrails. That said, there might not be a clear cut on the responsibilities (notably, robustness) between LLMs and the guardrails, and the two models shall collaborate to achieve a joint set of objectives. Nevertheless, for concrete applications, the requirements need to be precisely defined, together with their corresponding metrics, and a multi-disciplinary approach is called for. The mitigation of a given requirement (such as hallucinations, toxicity, fairness, biases, etc) is already non-trivial, as discussed in Section 3. The need to work with multiple requirements makes it worse, especially when some requirements can be conflicting. Such complexity requires a sophisticated solution design method to manage. In terms of the design of guardrails, while there might not be “one method that rules them all”, a plausible design of the guardrail is neural-symbolic, with learning agents and symbolic agents collaborating in processing both the inputs and the outputs of LLMs. There are multiple types of neural-symbolic agents (Lamb et al., 2021). However, the existing guardrail solutions such as Llama Guard (Inan et al., 2023), Nvidia NeMo (Rebedea et al., 2023), and Guardrails AI (Rajpal, 2023) use the simplest, loosely coupled ones. Given the complexity of the guardrails, it will be interesting to investigate other, more deeply coupled, neural-symbolic solution designs.
Like safety-critical software, a systematic process to cover the development cycle (ranging from specification, to design, implementation, integration, verification, validation, and production release) is required to carefully build the guardrails, as indicated in industrial standards such as ISO-26262 and DO-178B/C. The goal of this paper is to review the state-of-the-art (Section 2), present technical challenges on implementing individual requirements (Section 3), and then discuss several issues regarding the systematic design of a guardrail for a specific application context (Section 4).
2 Existing Implementation Solutions
This section reviews three existing implementation solutions for guardrails [^1] There are other guardrails available in the market, such as Open AI’s solution, Microsoft Azure AI Content Safety, Google Guardrails for Generative AI. However, they are either not open-sourced or lack details and contents for reproduction. Our discussion is limited to the three guardrails that are open-source and have been successfully replicated in our experiments., and discusses their pros and cons.
$Llama\ Guard$ (Inan et al., 2023), developed by Meta on the Llama2-7b architecture, focuses on enhancing Human-AI conversation safety. It is a fine-tuned model that takes the input and output of the victim model as input and predicts their classification on a set of user-specified categories. Figure 1 shows its workflow. Due to the zero/few-shot abilities of LLMs, Llama Guard can be adapted–by defining the user-specified categories –to different taxonomies and sets of guidelines that meet requirements for different applications and users. This is a Type 1 neural-symbolic system (Lamb et al., 2021), i.e., typical deep learning methods where the input and output of a learning agent are symbolic. It lacks guaranteed reliability since the classification results depend on the LLM’s understanding of the categories and the model’s predictive accuracy.
Figure 1: Llama Guard Guardrail Workflow
$Nvidia\ NeMo$, described in (Rebedea et al., 2023), functions as an intermediary layer that enhances the control and safety of LLMs. It employs Colang, an executable programme language designed by Nvidia (2023), to establish constraints, in order to guide LLMs within set dialogical boundaries. When the customer’s input prompt comes, NeMo embeds the prompt as a vector, and then uses K-nearest neighbor (KNN) method to compare it with the stored vector-based user canonical forms, retrieving the embedding vectors that are ‘the most similar’ to the embedded input prompt. After that, Nemo starts the flow execution to generate output from the canonical form. During the flow execution process, the LLMs are used to generate a safe answer if requested by the Colang program. The process is presented in Figure 2. Building on the above customizable workflow, NeMo also includes a set of pre-implemented moderations dedicated to e.g., fact-checking, hallucination prevention in responses, and content moderation. NeMo is also a Type-1 neural-symbolic system, with its effectiveness closely tied to the performance of the KNN method.
Figure 2: Nvidia NeMo Guardrails Workflow
$Guardrails\ AI$ enables the user to add structure, type and quality guarantees to the outputs of LLMs (Rajpal, 2023). It operates in three steps: 1) defining the “RAIL” spec, 2) initializing the “guard”, and 3) wrapping the LLMs. In the first step, Guardrails AI defines a set of RAIL specifications, which are used to describe the return format limitations. This information is required to be written in a specific XML format, facilitating subsequent output checks, e.g., structure and types. The second step involves activating the defined spec as a guard. For applications that require categorized processing, such as toxicity checks, additional classifier models can be introduced to categorize the input and output text. The third step is triggered when the guard detects an error. Here, the Guardrails AI can automatically generate a corrective prompt, pursuing the LLMs to regenerate the correct answer. The output is then re-checked to ensure it meets the specified requirements. Currently, the methods based on Guardrails AI are only applicable for text-level checks and cannot be used in multimodal scenarios involving images or audio. Unlike the previous two methods, Guardrail AI is a Type-2 neural-symbolic system, which consists of a backbone symbolic algorithm supported by learning algorithms (in this case, those additional classifier models).
Figure 3: Guardrails AI Workflow
Nevertheless, these solutions only provide the basic infrastructure (language for rule description, example workflow), without comprehensive studies on if and how such infrastructure can be utilized to implement a satisfactory guardrail. Research is needed to understand detailed issues regarding the infrastructures, including their capability (in dealing with, e.g., configuration redundancy and conversational capability limitations), generalization (in dealing with unforeseen scenarios), and expressivity (of enabling suitable interactions of symbolic and learning components). More importantly, a systematic approach of building guardrails based on the infrastructures is called for.
3 Technical Challenges of Implementing Individual Requirements
This section will review the technical challenges of implementing individual requirements, highlighting the intriguing complexity of dealing with a requirement. We consider four categories of requirements that might be requested in a specific context or application. Table 1 provides a summarisation of existing representative works. For every category of requirements, it classifies techniques into three groups. For vulnerability detection, the victim LLMs are typically treated as a blackbox, and thus they can be either with or without guardrails. Protection via LLMs enhancement includes techniques that tune the weights of LLMs. In contrast, For protection via I/O engineering, we consider any techniques that work on input and output, e.g., prompt engineering and output filter.
Vulnerability Detection Protection via LLMs Enhancement Protection via I/O Engineering
Free from Unintended Response
(Kang et al., 2023) (Wei et al., 2023)
(Shen et al., 2023)(Deng et al., 2023)
(Yong et al., 2023)(Vega et al., 2023)
(Zhang & Ippolito, 2023)(Albert, 2024)
(Li et al., 2018)(Liu et al., 2020)
(Miyato et al., 2016)(Ganguli et al., 2022)
(Touvron et al., 2023)(Perez et al., 2022)
(Askell et al., 2021)(Nakano et al., 2021)
(Jain et al., 2023)(Kumar et al., 2023)
(Robey et al., 2023)(Kim et al., 2023)
(Rajpal, 2023)(Inan et al., 2023)
(Rebedea et al., 2023)
Fairness
(Koh et al., 2023)(Motoki et al., 2023)
(Limisiewicz et al., 2023)(Badyal et al., 2023)
(Yeh et al., 2023)(Shaikh et al., 2022)
(Ranaldi et al., 2023)(Limisiewicz et al., 2023)
(Xie & Lukasiewicz, 2023)(Ernst et al., 2023)
(Ungless et al., 2022)(Ramezani & Xu, 2023)
(Huang et al., 2023a)
(Tao et al., 2023)(Oba et al., 2023)
(Dwivedi et al., 2023)
Privacy
(Zou et al., 2023) (Wang et al., 2023b)
(Li et al., 2023b)(Huang et al., 2022)
(Li et al., 2023a) (Lukas et al., 2023)
(Wang et al., 2024a)(Mireshghallah et al., 2023)
(Zanella-Béguelin et al., 2020)(Shi et al., 2022)
(Igamberdiev & Habernal, 2023)(Yu et al., 2022)
(Mireshghallah et al., 2022)(Xiao et al., 2023)
(Ozdayi et al., 2023)
(Li et al., 2023c)
(Duan et al., 2023)
Hallucination
(Ji et al., 2023)(Manakul et al., 2023)
(Bang et al., 2023)(Chen & Shu, 2023)
(Xu et al., 2024)(Huang et al., 2023b)
(Chern et al., 2023) (Cohen et al., 2023)
(Meng et al., 2022b)(Chuang et al., 2023)
(Meng et al., 2022a) (Bayat et al., 2023)
(Wang et al., 2024b)(Elaraby et al., 2023)
(Liang et al., 2024)(Razumovskaia et al., 2023)
(Press et al., 2022)(Gao et al., 2023)
(Pinter & Elhadad, 2023)(He et al., 2022)
(Zhao et al., 2023)(Ram et al., 2023)
(Dhuliawala et al., 2023)(Wang et al., 2023a)
Table 1: Literature on detecting and mitigating individual risks.
3.1 Free from Unintended Response
Recent studies have highlighted a growing concern about the ability of LLMs like ChatGPT to generate toxic contents, even with guardrails in place (Burgess, 2023; Christian, 2023; Zou et al., 2023). Most research uses prompt engineering methods to cause LLMs to create unintended content, a process often referred to as “jailbreaking”.
Vulnerability Detection Kang et al. (2023); Wei et al. (2023); Shen et al. (2023); Deng et al. (2023) have demonstrated that the LLMs can be manipulated to produce malicious contents using specific prompts. In addition, Kang et al. (2023) used text-davinci-003 prompt, Wei et al. (2023) explored failure models, Shen et al. (2023) employed “DAN” (Do Anything now”), Zou et al. (2023) introduced automated prompt generation based on gradient, Deng et al. (2023) proposed a balanced way by combining manual and automatic prompt generation together, and Vega et al. (2023) created few-shot priming attack and forced the LLMs to start generating from the middle of a sentence. Zhang & Ippolito (2023) evaluated the effectiveness of these prompt manipulation attacks. Beyond that, Yong et al. (2023) bypassed GPT-4’s safeguard by translating the English inputs into low-source languages. During our tests, we observed that certain vulnerabilities in LLMs that were previously known have been addressed, possibly due to the updates made by developers to enhance security measures. Nonetheless, a considerable number of individuals referred to as “jailbreakers” remain capable of effectively deceiving ChatGPT, which are tested in the publicly accessible project (Albert, 2024), as demonstrated in Appendix B.
Protection via LLMs Enhancement The LLMs can be enhanced by inherent safety training technologies. It can be achieved via the augmentation of training data by adding adversarial examples (Li et al., 2018; Ganguli et al., 2022; Perez et al., 2022; Mozes et al., 2023). Moreover, various efforts have been made to enhance safety during the RLHF process. Touvron et al. (2023) proposed to incorporate a safety reward into the RLHF process to prevent harmful outputs. Askell et al. (2021) improved the RLHF process by implementing context distillation in the training dataset. In the context of LLMs, Nakano et al. (2021) used the Reject Sampling mechanism to select the least harmful responses, thereby shaping the training dataset for RLHF. The robustness of language models can also be improved by modifying the training loss functions (Liu et al., 2020; Miyato et al., 2016). However, these adaptions are ineffective for the LLMs due to the catastrophic forgetting in the training process (Jain et al., 2023). Furthermore, these approaches require retraining of the LLMs to defend against the attacks, which can be unsuitable due to high-cost and closed-source nature.
Protection via I/O Engineering While detection and model enhancement are crucial, they alone are insufficient to safeguard against the evolving nature of threats, especially in the scenario where the model is not open. Consequently, several I/O engineering approaches that work on the input/output prompts have emerged. Jain et al. (2023) explored various defense technologies, including preprocessing and rephrasing input prompts. Kumar et al. (2023) used a safety filter on input prompts for certified robustness. Robey et al. (2023) introduced randomized smoothing technology to defend against such attacks by modifying input prompts and using majority voting for detection. Additionally, guardrail tools such as Guardrails AI and Nemo also offer detection and protection functions for harmful and toxic outputs.
Our Perspective As Tramer et al. (2020) have pointed out, while the defenses are effective against certain attacks, they remain vulnerable to stronger ones. This could turn into a continuous and infinite cycle of attacks and defenses. Consequently, a more robust solution is required, ideally offering provable guarantees to confirm the LLMs’ robustness against all adversarial attacks within a permissible perturbation limit. Toward this goal, we notice that existing guardrails seldom consider providing such guarantees. First and foremost, it is necessary to develop metrics for toxicity and other criteria to address unintended responses. In terms of these metrics, rather than relying on purely empirical measures which may improve the performance but cannot lead to guarantees, we can consider certified robustness bounds, either statistical bound (Cohen et al., 2019; Zhao et al., 2022) or deterministic bound (Huang et al., 2017; Sun & Ruan, 2023), as scores to measure the guardrail performance. Additionally, we can also incorporate the metrics (or the bounds) into the training process of the LLMs for improvement, or use it in the fine-tuning process.
3.2 Fairness
Fairness in LLMs has been studied from different angles, such as gender bias (Malik, 2023; Sun et al., 2023; Ovalle et al., 2023), cultural bias (Tao et al., 2023; Gupta et al., 2023), dataset bias (Sheppard et al., 2023), and social bias (Sheng et al., 2023; Manerba et al., 2023; Tang et al., 2023; Gonçalves & Strubell, 2023; Nagireddy et al., 2023; Bi et al., 2023). Understanding and addressing biases in LLMs requires solid theoretical frameworks and comprehensive analysis. Gallegos et al. (2023) provided a comprehensive overview of social biases and fairness in natural language processing, offering a framework for identifying and categorizing different types of harms, intuitive taxonomies for bias evaluation metrics and datasets, and a guide for mitigations.
Vulnerability Detection Badyal et al. (2023) purposefully incorporated biases into the responses of LLMs to craft distinct personas for use in interactive media. Koh et al. (2023) focused on identifying and quantifying instances of social bias in models like ChatGPT, especially in sensitive applications such as job and college admissions screening. Limisiewicz et al. (2023) proposed a novel method for detecting gender bias in language models. Motoki et al. (2023) examined the presence of political bias in ChatGPT, focusing on aspects such as race, gender, religion, and political orientation. Additionally, they explored the role of randomness in responses, by collecting multiple answers to the same questions, which enables a more robust analysis of potential biases. Yeh et al. (2023) examined the bias of LLMs by controlling the input, highlighting that LLMs can still produce biased responses despite the progress in bias reduction. Shaikh et al. (2022) designed a Bias Index to quantify and address biases inherent in LLMs including GPT-4. It has also been observed that the biased response can be generated inadvertently, sometimes in the form of seemingly harmless jokes (Zhou & Sanfilippo, 2023) (demonstrated in Appendix B). Such instances may not be sufficiently addressed by existing guardrail systems.
Protection via LLMs Enhancement Many studies have concentrated on reducing bias through model adaption approaches. Limisiewicz et al. (2023) provided a bias mitigating method, DAMA, that can reduce bias while maintaining model performance on downstream tasks. Ranaldi et al. (2023) investigated the bias in CtB-LLMs and demonstrate the effectiveness of debiasing techniques. They find that bias is not solely dependent on the number of parameters but also on factors like perplexity, and that techniques like debiasing of OPT using LoRA can significantly reduce bias. Ungless et al. (2022) demonstrated that the Stereotype Content Model, which posits that minority groups are often perceived as cold or incompetent, applies to contextualized word embeddings and presents a successful fine-tuning method to reduce such biases. Moreover, Ernst et al. (2023) proposed a novel adversarial learning debiasing method, applied during the pre-training of LLMs. Ramezani & Xu (2023) mitigated cultural bias through fine-tuning models on culturally relevant data.
Protection via I/O Engineering In addition to fine-tuning methods, several studies exploring the control of input and output. Huang et al. (2023a) suggested to use purposely designed code generation templates to mitigate the bias in code generation tasks. Tao et al. (2023) found that cultural prompting is a simple and effective method to reduce cultural bias in the latest LLMs, although it may be ineffective or even exacerbate bias in some countries. Oba et al. (2023) proposed a method to address gender bias that does not require access to model parameters. It shows that text-based preambles generated from manually designed templates can effectively suppress gender biases with minimal adverse effects on downstream task performance. Dwivedi et al. (2023) guided LLMs to generate more equitable content by employing an innovative approach of prompt engineering and in-context learning, significantly reducing gender bias, especially in traditionally problematic.
Our Perspective To effectively mitigate bias, it’s crucial to develop guardrails through a comprehensive approach that intertwines various strategies. This begins with meticulously monitoring and filtering training data to ensure it is diverse and devoid of biased or discriminatory content. The essence of this step lies in either removing biased data or enriching the dataset with more inclusive and varied information. Alongside this, algorithmic adjustments are necessary, which involve fine-tuning the model’s parameters to prevent the overemphasis of certain patterns that could lead to biased outcomes. Incorporating bias detection tools is another pivotal aspect. These tools are designed to scrutinize the model’s outputs, identifying and flagging potentially biased content for human review and correction. We believe that a key to the long-term efficacy of these guardrails is the adoption of a continuous learning approach. This involves regularly updating the model with new data, insights, and feedback and adapting to evolving societal norms and values. This dynamic process ensures that the guardrails against bias remain robust and relevant. Moreover, the above issues can and should be addressed with a multidisciplinary team, as discussed in Section 4.2. Also, similar to the discussion in Section 3.1, we believe in principled methods to evaluate fairness when the definitions are clearly settled. It is however expected that the definition will be distribution-based, rather than point-based as unintended responses, which need to estimate posterior distributions and to measure the distance between two distributions.
3.3 Privacy and Copyright
Legislations such as the EU AI Act, General Data Protection Regulation (GDPR), and California Consumer Privacy Act (CCPA) have established rigorous standards for data sharing and retention. These frameworks mandate strict compliance with data protection and privacy guidelines. Privacy-related research focuses on the risks of either leaking training data or the trained model. The former includes the attacks and defense on e.g., determining if a data point is within the training dataset (Shokri et al., 2017), reconstructing a training data point from a subset of the features (Zhang et al., 2020), or reconstructing some of the training data (Balle et al., 2022). The latter infers information from the model, see e.g., (Wang et al., 2021). In the following, we focus on the privacy on the training data.
Vulnerability Detection LLMs face the challenges in releasing the personal identifiable information (PII) (Li et al., 2023b, a; Lukas et al., 2023; Huang et al., 2022; Wang et al., 2024a), highlighting the need for caution and robust data handling protocols. They are pre-trained on extensive textual datasets (Narayanan et al., 2021) and can inadvertently reveal sensitive information about data subjects (Plant et al., 2022). Specifically, Li et al. (2023b) considered the risks of leaking personal information in e.g., text completion task where the adversary attempts to recover private information by using tricky prompt as the prefix, and Wang et al. (2024a) used an aggregated score to evaluate the LLM’s privacy. Mireshghallah et al. (2023) also exhaustively tested the latest ChatGPT about their capability of keeping a secret.
Protection via LLMs Enhancement Numerous studies have focused on implementing privacy defense technologies to safeguard data and model privacy and counter privacy breaches, with the Differential Privacy (DP) based methods (Abadi et al., 2016) as the most studied. For general NLP models, Li et al. (2022) indicated that a direct application of DP-SGD (Abadi et al., 2016) may not achieve satisfactory performance, and suggests a few tricks. Igamberdiev & Habernal (2023) implemented a model for text rewriting along with Local Differential Privacy (LDP), both with and without pretraining. For LLMs, the focus has been on the integration of DP into the fine-tuning process (Yu et al., 2022; Shi et al., 2022; Mireshghallah et al., 2022). Other than DP-based methods which deal with general differential privacy, Xiao et al. (2023) considered contextual privacy, which measures the sensitivity of a piece of information upon the context, and injects domain-specific knowledge into the fine-tuning process.
Protection via I/O Engineering Ozdayi et al. (2023) proposed a method to prepend a trained prompt to the incoming prompt before passing them to the model, where the training of the prefix prompt is to minimise the extent of extractable memorized content in the model. Li et al. (2023c) and Duan et al. (2023) also proposed the prompt-tuning methodology that adhere to differential privacy principles.
Our Perspective Other than constructing privacy-preserving LLMs, watermarking techniques can play a more important role in LLMs, for not only privacy but also copyright protection. A typical watermarking mechanism (Kirchenbauer et al., 2023) embedded watermarks into the output of LLMs by selecting a randomized set of “green” tokens before a word is generated, and then softly promoting the use of green tokens during sampling. So, as long as we know the list of green tokens, it is easy to determine if an output is watermarked or not. We can also use the watermarks to track the point of origin or the owner of watermarked text for copyright purposes, and this has been applied to protect the copyright of generated prompts (Yao et al., 2023). We believe in an agreed watermarking mechanism between the data owners and the LLMs developers, such that the users embed a personalized watermark into their documents or texts when they deem them private or with copyright, and the LLMs developers will not use watermarked data for their training. More importantly, the LLMs developers should take the responsibility of enabling (1) an automatic verification to determine if a user-provided, watermarked text is within the training data, and (2) model unlearning (Nguyen et al., 2022), which allows the removal of users’ personally owned texts from training data.
3.4 Hallucinations and Uncertainty
LLMs have a notable inclination to generate hallucinations (Ji et al., 2023; Bang et al., 2023), leading to contents that deviate from real-world facts or user inputs. The hallucinations in conditional text generation are closely tied to high model uncertainty (Huang et al., 2023b). The absence of uncertainty measures for LLMs significantly hampers the reliability of information generated by LLMs.
Vulnerability Detection Chen & Shu (2023) first identified the challenges in detecting the misinformation in ChatGPT, resulting in a growing number of research to explore the factual hallucination that is inconsistent with real-world facts. Chern et al. (2023) proposed a cohesive framework by utilizing a range of external tools for gathering evidence to identify factual inaccuracies. Some methods aim to detect hallucinations without relying on external sources by focusing on the model’s uncertainty in generating factual content. Manakul et al. (2023) proposed to identify hallucinations by generating multiple responses and evaluating the consistency of factual statements. Apart from evaluating uncertainty through the self-consistency of multiple generations from a single LLM, one can adopt a multi-agent approach by including additional LLMs (Cohen et al., 2023). Worsely, Xu et al. (2024), claim that LLMs cannot completely eliminate hallucinations. They define a formal world where hallucination is characterized as inconsistencies between computable LLMs and a computable ground truth function.
Protection via LLMs Enhancement Meng et al. (2022b) proposed to mitigating data-related hallucinations in LLMs by increasing the amount of factual data during the pre-training phase, and this proposal was later refined by Meng et al. (2022a). Modifying the training dataset can partially reduce the model’s knowledge gap effectively. Besides, Liang et al. (2024) developed an automated hallucination annotation tool, DreamCatcher, and proposing a Reinforcement Learning from Knowledge Feedback training framework, effectively improving their performance in tasks related to factuality and honesty. Wang et al. (2024b) introduced the ReCaption framework, which combines rewriting captions using ChatGPT with fine-tuning large vision-language models, successfully reducing fine-grained object hallucinations in LVLMs More related works can be found in (Tonmoy et al., 2024).
Protection via I/O Engineering Apart from the refining methods, Pinter & Elhadad (2023) found that these methods might pose potential risks when trying to combat LLMs hallucinations. They recommend using retrieval-augmented methods, which seek to add external knowledge acquired from retrieval directly to the LLMs’ prompt (He et al., 2022; Press et al., 2022; Ram et al., 2023). Based on the Chain-of-Thought technology, Dhuliawala et al. (2023) introduced the “Chain-of-Verification” method to effectively reduce the generation of inaccurate information in LLMs. Wang et al. (2023a) then proposed a faithful knowledge distillation method that significantly enhances the credibility and accuracy of LLMs. Zhao et al. (2023) proposed a Verify-and-Edit framework based on GPT-3, which enhances the factual accuracy of predictions in open-domain question-answering tasks. Additionally, Gao et al. (2023) pioneered the “Retrofit Attribution using Research and Revision” system, which improves the outputs by automatically attributing and post-editing generated text to correct inaccuracies.
Our Perspective As suggested earlier, uncertainty can be utilized to deal with hallucinations. The primary challenges of LLMs uncertainty stem from the critical roles of meaning and form in language. This relates to what linguists and philosophers refer to as a sentence’s semantic content and its syntactic or lexical structure. Foundation models primarily produce token-likelihoods, indicating lexical confidence. However, in most applications, it is the meanings that are of paramount importance. Kuhn et al. (2022) presented the concept of semantic entropy, an entropy that integrates linguistic invariances brought about by the same meaning. The fundamental method involves a semantic equivalence relation to express that two sentences have the same meaning.
In addition, we need to consider the uncertainty of the measurements of LLM. For example, for the assessment of toxicity levels, there are quantitative methods like tracking the frequency of toxic words or using sentiment analysis scores, and qualitative approaches such as evaluations by experts. It is crucial to verify that these metrics are consistent and applicable across a variety of contexts and content types. We also highlight the need to account for the inherent uncertainty of LLMs, an aspect not sufficiently addressed in previous guardrail designs. Incorporating uncertainty measurements such as conformal predictions (Shafer & Vovk, 2008) could enhance the evaluation of fairness, creativity, and privacy of LLMs in generating questions by considering the uncertainty level and all possible responses.
4 Challenges on Designing Guardrails
Based on the discussions about tackling individual requirements as discussed in Section 3, this section advocates the building of a guardrail by considering multiple requirements in a systematic way. We discuss four topics: conflicting requirements (Section 4.1), multidisciplinary approach (Section 4.2), implementation strategy (Section 4.3), and rigorous engineering process (Section 4.4).
4.1 Conflicting Requirements
This section discusses the tension between safety and intelligence, as an example for the conflicting requirements. Conflicting requirements are typical, including e.g., fairness and privacy (Xiang, 2022), privacy and robustness (Song et al., 2019), and robustness and fairness (Bassi et al., 2024).
The integration of guardrails with LLMs may lead to a discernible conservative shift in the generation of responses to open-ended text-generation questions (Röttger et al., 2023). The shift has been witnessed in ChatGPT over time. Chen et al. (2023) documented a notable change in ChatGPT’s performance between March and June 2023. Specifically, when responding to sensitive queries, the model’s character count decreased significantly, plummeting from an excess of 600 characters to approximately 140. Additionally, in the context of opinion-based questions and answers surveys, the model is more inclined to abstain from responding.
Given the brevity and conservativeness of responses generated by ChatGPT, it raises the question: How can exploratory depth be maintained in responses, particularly for open-ended test generation tasks? Furthermore, does the application of guardrails constrain ChatGPT’s capacity to deliver more intuitive responses? On the other hand, Narayanan & Kapoor (2023) critically examined this paper, and emphasized the difference between an LLM’s capabilities and its behavior. Although capabilities typically remain constant, behavior can alter due to fine-tuning, which can be interpreted as the “uncertainty” challenges in LLMs. They suggest that changes in GPT-4’s performance are likely linked more to evaluation data and fine-tuning methods rather than a decline in its fundamental abilities. They also acknowledge that such behavioral drift poses a challenge in developing reliable chatbot products. The adoption of guardrail has also led to the model adopting a more succinct communication approach, thereby offering fewer details or electing for non-response in certain queries. The decision of “to do or not to do” can be a challenging task when designing the guardrail. While the easiest approach is to decline an answer to any sensitive questions, is it the most intelligent one? That is, we need to determine if the application of guardrail always has a positive impact on LLMs that is within our expectation.
Our Perspective Prior research has suggested to incorporate a creativity assessment mechanism into the guardrail development for LLMs. To measure the creativity capability of LLMs, Chakrabarty et al. (2023) employed the Consensual Assessment Technique (Amabile, 1982), a well-regarded approach in creativity evaluation, focusing on several key aspects: fluency, flexibility, originality, and elaboration, which collectively contribute to a comprehensive understanding of the LLMs’ creative output in storytelling. Narayanan & Kapoor (2023) showed that although some LLMs may demonstrate adeptness in specific aspects of creativity, there is a significant gap between their capabilities and human expertise when evaluated comprehensively.
4.2 Multidisciplinary Approach
While current LLMs guardrails include mechanisms to detect harmful contents, they still pose a risk of generating biased or misleading responses. It is reasonable to expect the future guardrails to integrate not only harm detections but also other mechanisms to deal with, e.g., ethics, fairness, and creativity. We have provided in the Introduction three categories of requirements to be considered for a guardrail. Moreover, LLMs may not be universally effective across all domains, and it has been a trend to consider domain-specific LLMs (Pal et al., 2023). In domain-specific scenarios, specialized rules may conflict with the general principles. For instance, in crime prevention, the use of certain terminologies that are generally perceived as harmful, such as ‘guns’ or ‘crime’, is predominant and should not be precluded. To this end, the concrete requirements for guardrails will be different across different LLMs, and research is needed to scientifically determine requirements. The above challenges (multiple categories, domain-specific, and potentially conflicting requirements) are compounded by the fact that many requirements, such as fairness and toxicity, are hard to be precisely defined, especially without a concrete context. The existing methods, such as the popular one that sets a threshold on predictive toxicity level (Perez et al., 2022), do not have valid justification and assurance.
Our Perspective Developing LLMs ethically involves adhering to principles such as fairness, accountability, and transparency. These principles ensure that LLMs do not perpetuate biases or cause unintended harm. The works by e.g., Sun et al. (2023) and Ovalle et al. (2023) provide insights into how these principles can be operationalized in the context of LLMs. Establishing community standards is vital for the responsible development of LLMs. These standards, derived from a consensus among stakeholders, including developers, users, and those impacted by AI, can guide the ethical development and deployment of LLMs. They ensure that LLMs are aligned with societal values and ethical norms, as discussed in broader AI ethics literature (ActiveFence, 2023). Moreover, the ethical development of LLMs is not a one-time effort but requires ongoing evaluation and refinement. This involves regular assessment of LLMs outputs, updating models to reflect changing societal norms, and incorporating feedback from diverse user groups to ensure that LLMs remain fair and unbiased.
Socio-technical theory (Trist & Bamforth, 1957), in which both ‘social’ and ‘technical’ aspects are brought together and treated as interdependent parts of a complex system, have been promoted (Filgueiras et al., 2023; Jr. et al., 2020) for machine learning to deal with properties related to human and societal values, including e.g., fairness (Dolata et al., 2022), biases (Schwartz et al., 2022), and ethics (Mbiazi et al., 2023). To manage the complexity, the whole system approach (Crabtree et al., 2011), which promotes an ongoing and dynamic way of working and enables local stakeholders to come together for an integrated solution, has been successfully working on healthcare systems (Brand et al., 2017). We believe, a multi-disciplinary group of experts will work out, and rightly justify and validate, the concrete requirements for a specific context, by applying the socio-technical theory and the whole system approach.
4.3 Neural-Symbolic Approach for Implementation
Existing guardrail frameworks such as those introduced in Section 2 employ a language (such as RAIL or Colang) to describe the behavior of a guardrail. A set of rules and guidelines are expressed with the language, such that each of them is applied independently. It is unclear if and how such a mechanism can be used to deal with more complex cases where the rules and guidelines have conflicts. As mentioned in Section 4.2, such complex cases are common in building guardrails. Moreover, it is unclear if they are sufficiently flexible, and capable of adapting, to semantic shifts over time and across different scenarios and datasets.
Our Perspective First, a principled approach is needed to resolve conflicts in requirements, as suggested in (van Lamsweerde et al., 1998) for requirement engineering, which is based on the combination of logic and decision theory. Second, a guardrail requires the cooperation of symbolic and learning-based methods. For example, we may expect that, the learning agents deal with the frequently-seen cases (where there are plenty of data) to improve the overall performance w.r.t. the above-mentioned requirements, and the symbolic agents take care of the rare cases (where there are few or no data) to improve the performance in dealing with corner cases in an interpretable way. Due to the complex conflict resolution methods, more closely-coupled neural-symbolic methods might be needed to deal with the tension between effective learning and sound reasoning, such as those Type-6 systems (Lamb et al., 2021) that can deal with true symbolic reasoning inside a neural engine, e.g., Pointer Networks (Vinyals et al., 2015).
4.4 Systems Development Life Cycle (SDLC)
The criticality of guardrails requires a careful engineering process to be applied, and for this, a revisit of the SDLC, which is a complex project management model to encompass guardrail creation from its initial idea to its finalized deployment and maintenance has the potential, and the V-model (Oppermann, 2023), which builds the relations of each development process with its testing activities, can be useful to ensure the quality of the final product.
Our Perspective Rigorous verification and testing will be needed (Huang et al., 2023b), which requires a comprehensive set of evaluation methods. For individual requirements, certification with statistical guarantees can be useful, such as the randomized smoothing (Cohen et al., 2019). For the evaluation of multiple, conflicting requirements, a combination of the Pareto front based evaluation methods for multiple requirements (Ngatchou et al., 2005) and the statistical certification for a single requirement is needed. Finally, attention should also be paid to understanding the theoretical limits of the evaluation methods, e.g., randomized smoothing causes a fairness problem (Mohapatra et al., 2021).
5 Conclusion
This position paper advocates for a systematic approach to building guardrails, beyond the current solutions which only offer the simplest mechanisms to describe rules and connect learning and symbolic components. Guardrails are highly complex due to their role of managing interactions between LLMs with humans. A systematic approach, supported by a multidisciplinary team, can fully consider and manage the complexity and provide assurance to the final product.
Impact Statement
This paper shares our views about how to build a responsible safeguarding mechanism for Large Language Models (LLMs), a generative AI technique. In this sense, it holds positive societal impacts. Nevertheless, to expose the problems, the paper also includes example questions and model outputs that may be perceived as offensive.
Appendix A Comparison of Llama Guard, Nemo and Guardrails AI
We build the qualitative analysis dimensions based on the workflow of the guardrails (refer to Figure 1, Figure 2 and Figure 3), as shown in Table 2. The first factor to take into account is the capability of customizing rules for guardrails. Customized rules are considered into two dimensions, where Monitoring rules refers to the ability to allow users to customize the functions performed by guardrails, and Enforcement rules denotes the capacity to compel the production of predefined content upon detection of content. It is noted here that LLama Guard only classifies the output text, but does not enforce the output. Multi-modal support considers whether the input-output properties of the guardrail support multi-modality. Guardrails AI can only support text-based checks. In terms of the Output check, Nemo’s output follows the flow execution of the Colang program, but there is no further validation if it imported GPT’s generation. Scalability support demonstrates whether the guardrail framework is applicable to the specific LLM. Llama Guard checks users’ input and LLM’s output, and does not interact directly with LLMs, so it is not considered for this dimension. Nvidia NeMo is only available with ChatGPT and Guardrails AI provides better scalability support.
Llama Guard Nvidia NeMo Guardrails AI
Monitoring rules
Enforcement rules
Multi-modal support
Output check
Scalability support
Table 2: Compared Results of Guardrail Frameworks under Qualitative Analysis Dimensions
Appendix B Demonstration of the current challenges in ChatGPT
In this section, we showcase the negative aspects of ChatGPT’s responses, as depicted in Figure 4. These aspects include unintended responses, biases, privacy breaches, and hallucinations. Additionally, we demonstrate the challenges faced by current guardrailed chatbots when it comes to refusing responses and delivering overly cautious responses in Figure 5.
In Figure 4(a), when we change the input prompt to a “Hypothetical response”, ChatGPT provides a step-by-step guide on an illegal act, such as hotwiring a car, raising significant safety concerns. In the example illustrated in Figure 4(b), an unfair response may inadvertently come across as a joke, assuming that fairness is widely accepted or understood without further explanation. This oversight can have negative repercussions on communities, especially children, as it perpetuates harmful biases without adequate context or explanation. Regarding privacy leakage, we demonstrate an example in Figure 4(c), revealing that ChatGPT is unable to keep a secret within the conversation, even when we mention that the message will be shared with all attendees. In 4(d), we observe that when ChatGPT is asked to provide references, some information in the references can be inaccurate, raising concerns about the reliability in the scientific information.
In the context of opinion-based questions and answers surveys, the model is more inclined to abstain from responding, as demonstrated in Figure 5. As we can see in the example presented in Figure 5 (b), ChatGPT 4 tends to decline to answer potentially sensitive questions, even abstaining from delivering positive responses.
(a) Jailbreaking
(b) Fairness
(c) Privacy Leakage
(d) Hallucination
Figure 4: Harmful Response
(a) Refuse Response
(b) Conservative
Figure 5: Safer or Intelligence? How to Respond
Appendix C Evaluate current attack method
In this section, we showcase that ChatGPT 3.5 and 4 have successfully addressed and resolved certain state-of-the-art attack methods. Representative examples are shown in Figures 6(a)-6(d).
(a) Instruction-following attack (Kang et al., 2023)
(b) Jailbroken (Wei et al., 2023)
(c) Jailbroken: Do Anything Now (Shen et al., 2023)
(d) Hypothetical response (Wei et al., 2023)
Figure 6: Guarded Example of Attacks on ChatGPT